OpenStudy (anonymous):

What is the length of one side of a square with a 16-centimeter diagonal?

OpenStudy (anonymous):

√12 cm 8√2 cm √16 cm

OpenStudy (saifoo.khan):

8root2

OpenStudy (anonymous):

8√2 cm

OpenStudy (anonymous):

a^2 + b^2 = 16^2, a=b a = b = 8sqrt(2)